OpenStudy (isbella):

1)test reivew:Let F⃗ (x,y,z)=6x2i⃗ −sin(xz)(i⃗ +k⃗ ). Calulate the divergence: div F⃗ (x,y,z)=

OpenStudy (unklerhaukus):

\[\boldsymbol{\vec F}(x,y,z) = 6x^2\boldsymbol{\hat\imath}-\sin(xz)(\boldsymbol{\hat\imath}+\boldsymbol{\hat k})\] \[\text{div}\boldsymbol{\vec F}(x,y,z) =\nabla\cdot\boldsymbol{\vec F}\\\qquad\qquad\qquad=\frac{\partial F_x}{\partial x}+\frac{\partial F_y}{\partial y}+\frac{\partial F_z}{\partial z}\]

OpenStudy (unklerhaukus):

First resolve the function into its components \[\boldsymbol{\vec F}(x,y,z)=F_x\boldsymbol{\hat\imath}+F_y\boldsymbol{\hat\jmath}+F_z\boldsymbol{\hat k}\]

OpenStudy (unklerhaukus):

Then take the respective partial derivatives of the components and add them together